MEASUREMENT Interconverting compound SI units Convert the following measurement. 3 mg 7.47 x 10 X
Answer #1

Solution:

Here dL means decilitre and mg means milligram.

1.00 g / L = 100 mg / dL

1.00 mg / dL = 0.01 g / L

where as 7.47*10^3 mg / dL = 0.01 * 7.47*10^3 g / L

                                                  = 7.47*10^1 g / L

                                                = 74.7 g / L

A:- 74.7 g / L